# Hidden-Role Games: Equilibrium Concepts and Computation

### v4

@@ Abstract
[…]
@@ 1.1 Main Modeling Contributions
[…]
@@ 1.2 Main Computational Contributions
[…]
@@ 1.4 Examples
Consider a $n$-player version of matching pennies (with $n>2$), which we denote as $\textsf{{MP}}(n)$. One player is chosen at random to be the adversary (team MIN). All $n$ players then simultaneously choose bits $b_{i}\in\{0,1\}$. Team MAX wins (gets utility $1$) if and only if all $n$ bits match; else, team MAX loses (gets utility 00).
With no communication, the value of this game is $1/2^{n-1}$: it is an equilibrium for everyone to play uniformly at random. Public communication does not help, because, conditioned on the public transcript, bits chosen by players must be mutually independent. Thus, the adversary can do the following: pretend to be on team MAX, wait for all communication to finish, and then play 00 if the string of all ones is more conditionally likely than the string of all $1$s, and vice-versa.
With private communication, however, the value becomes $1/(n+1)$. Intuitively, the MAX-team should attempt to guess who the MIN-player is, and then privately discuss among the remaining $n-1$ players what bit to play. We defer formal proofs of the above game values to Section 5, because they rely on results in Section 4.1.
@@ 2 Preliminaries
    A set of $n$ players, identified with the integers $[n]=\{1,\dots,n\}$, and an extra player C denoting the chance player modeling stochasticity.
@@
    For each player $i\in[n]$, a utility function $u_{i}:Z\to\mathbb{R}$.
@@
and the TMEs are the strategy profiles $x$ that achieve the maximum value. Notice that the TME problem is nonconvex, since the objective function $u$ is nonlinear as a function of $x$ and $y$. As such, the minimax theorem does not apply, and swapping the teams may not preserve the solution. Computing an (approximate) TME is $\Sigma_{2}^{\P}$-complete in extensive-form games (Zhang et al., 2023a).
@@ 3 Equilibrium Concepts for Hidden-Role Games
    The utility of a player $i$ is defined completely by its team: there is a $u:Z\to\mathbb{R}$ for which $u_{i}(z)=u(z)$ if player $i$ is on team MAX at node $z$, and $-u(z)$ otherwise. [^10] While at a first look this condition is similar to the one in ATGs, we remark that in this case the number of players in a team depends on the roles assigned at the start. The same considerations as Footnote 8 on the zero-sum rescaling of the utilities hold.
@@ 3.3 Equilibrium Notions
[…]
@@ 4.1 Computing Private-Communication Equilibria
[…]
@@ 4.2 Computing No/Public-Communication Equilibria
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 5 Worked Example
For the public-communication value, observe that, conditioned on the transcript, the bits chosen by the players must be mutually independent of each other. Thus, the adversary can do the following: pretend to be on team MAX, wait for all communication to finish, and then play 00 if the string of all ones is more conditionally likely than the string of all $1$s, and vice-versa [^13] In general, computing the conditional probabilities could take exponential time, but when defining the notion of value here, we are assuming that players have unbounded computational resources. This argument not work for computationally-bounded adversaries. Indeed, if the adversary were computationally bounded, MAX would be able to use cryptography to build private communication channels and thus implement a mediator, allowing our main positive result Theorem 4.3 to apply..
@@
(Lower bound) The players simulate a mediator using multi-party computation (see Theorems 4.3 and 4.2). Consider the following strategy for the mediator. Sample a string $b\in\{0,1\}^{n}$ uniformly at random from the set of $22n+$ 2 strings that has at most one mismatched bit. Recommend to each player $i$ that they play $b_{i}$.
@@
    Exactly one other bit of $b$ is $1$ ($n-1$ ways).
@@ 6.1 The Price of Hidden Roles
Let $D\in\Delta(\{{\color[rgb]{0.12109375,0.46484375,0.70703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.12109375,0.46484375,0.70703125}\textbf{{{MAX}}}},{\color[rgb]{1,0.49609375,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.49609375,0.0546875}\textbf{{{MIN}}}}\}^{n})$ be any distribution of teams assignments. Let $\mathcal{G}_{n,D}$ be the class of all hidden-role games with $n$ players and team assignment distribution $D$. Then the price of hidden roles of $\mathcal{G}_{n,D}$ is exactly the largest probability assigned to any team by $D$, that is,
@@
(Upper bound) Team MAX assumes that the true MAX-team is exactly the team $t^{*}$. Then MAX gets utility at most a factor of $1/p^{*}$ worse than the TME value of $\textup{{PublicTeam}}(\Gamma)$: if the assumption is correct, then MAX gets the TME value; if the assumption is incorrect, MAX gets value at least 00 thanks to the condition on MAX’s utilities in Definition 6.2.
(Lower bound) Consider the following game $\Gamma$. Nature first selects a team assignment $t\sim D$ and each player privately observes its team assignment. Then, all players are simultaneously asked to announce what they believe the true team assignment is. The MAX-team wins if every MAX-player announces the true team assignment. If MAX wins, MAX gets utility $1$; otherwise MAX gets utility 00.
@@
Let $\mathcal{G}_{n,k}$ be the class of all hidden-role games where the number of players and adversaries are always exactly $n$ and $k$ respectively. The price of hidden roles in $\mathcal{G}_{n,k}$ is exactly $\binom{n}{k}$.
@@ 7 Experimental Evaluation: Avalon
No special roles (Resistance) 3 / 10 $=$ 0.3000* 1 / 3 $\approx$ 0.3333*
Merlin 2 / 3 $\approx$ 0.6667* 3 / 4 $=$ 0.7500*
@@
A game of Avalon proceeds, generically speaking, as follows. There are $n$ players, $\left\lceil n/3\right\rceil$ of which are randomly assigned to team MIN and the rest to team MAX. Team MIN is informed. Some special rules allow players observe further information; for example, Merlin is a MAX-player who observes the identity of the players on team MIN, except the MIN-player Mordred, and the MAX-player Percival knows Merlin and Morgana (who is on team MIN), but does not know which is which. The game proceeds in five rounds. In each round, a leader publicly selects a certain number of people (defined as a function of the number of players and current round number) to go on a mission. Players then publicly vote on whether to accept the leader’s choice. If a strict majority vote to accept, the mission begins; otherwise, leadership goes to the player to the left. If four votes pass with no mission selected, there is no vote on the fifth mission (it automatically gets accepted). If a MIN-player is sent on a mission, they have the chance to fail the mission. The goal of MAX is to have three missions pass. If Merlin is present, MIN also wins by correctly guessing the identity of Merlin at the end of the game. Avalon is therefore parameterized by the number of players and the presence of the extra roles Merlin, Mordred, Percival, and Morgana.
@@
    A parallelized version of the PCFR+ algorithm (Farina et al., 2021), a scalable no-regret learning algorithm. PCFR+ was able to find an approximate equilibrium with exploitability $<10^{-3}$ in less than 10 minutes in the largest game instance, and was able to complete 10,000 iterations in under two hours for each game.
@@ Acknowledgements
The work of Prof. Sandholm’s group is supported by the National Science Foundation under grants RI2312342 and RI-1901403, and by the ARO under award W911NF2210266. Nicola Gatti is supported by FAIR (Future Artificial Intelligence Research) project, funded by the NextGenerationEU program within the PNRR-PE-AI scheme (M4C2, Investment 1.3, Line on Artificial Intelligence). We thank Justin Raizes for useful discussions on multi-party computation.
@@ B.1 Secure MPC
In secure MPC, there is a (possibly randomized) function $f:X^{n}\to X^{n}$ defined by a circuit with $N$ nodes. A subset $K\subset[n]$ of size $<n/2$ has been corrupted. Each honest player $i\in[n]\setminus K$ holds an input $x_{i}\in X$. The goal is to design a randomized messaging protocol, with private communication, such that, regardless of what the corrupted players do, there exist inputs $\{x_{j}:j\in K\}$ such that:
@@
    (Privacy) No subset of $<n/2$ players can learn any information except their own output $y_{i}$. That is, the players cannot infer any extra information from analyzing the transcripts of the message protocol than what they already know.
    This can be intuitively modeled as follows. If any minority of colluded players were to analyze the empirical distributions of transcripts of the protocol for any input-output tuple, then such distribution would be fully explained in terms of their input-outputs, leaving no conditional dependence on other players’ input-outputs. Formally, for any such subset $K\subseteq[n]$ of size $<n/2$, there exists a randomized algorithm ${\sf Sim}_{K}$ that takes the inputs and outputs of the players in set $K$, and reconstructs transcripts $T$, such that for all $x\in X^{n}$, we have
    $\sum_{T}\absolutevalue{\Pr[{\sf Sim}_{K}(x_{K},y_{K})=T]-\Pr[{\sf View}_{K}(x)=T]}\leq 2^{-\lambda}$ (4)
@@ B.2 Verifiable Secret Sharing
[…]
@@
    (Reconstructibility) Any subset of $>n/2$ players can recover the secret fully, even if the remaining players are adversarial. That is, there exists a function ${\sf Reconstruct}:X^{n}\to X$ such that, where $(x_{1},\dots,x_{n})\leftarrow{\sf Share}(x)$, we have ${\sf Reconstruct}(x_{1}^{\prime},\dots,x_{n}^{\prime})=x$ so long as $x_{i}^{\prime}=x_{i}$ for $>n/2$ players $i$.
@@
    (Privacy) No subset of $<n/2$ players can learn any information about the secret $x$. That is, for any such subset $K$ and any secret $x$, there exists a distribution ${\sf Sim}_{K}\in\Delta(X^{n})$ such that
@@
[…]

### v5

@@ Abstract
In this paper, we study the class of games known as hidden-role games in which players are assigned privately to teams and are faced with the challenge of recognizing and cooperating with teammates. This model includes both popular recreational games such as the Mafia/Werewolf family and The Resistance (Avalon) and many real-world settings, such as distributed systems where nodes need to work together to accomplish a goal in the face of possible corruptions. There has been little to no formal mathematical grounding of such settings in the literature, and it was previously not even clear what the right solution concepts (notions of equilibria) should be. A suitable notion of equilibrium should take into account the communication channels available to the players (e.g., can they communicate? Can they communicate in private?). Defining such suitable notions turns out to be a nontrivial task with several surprising consequences. In this paper, we provide the first rigorous definition of equilibrium for hidden-role games, which overcomes serious limitations of other solution concepts not designed for hidden-role games. We then show that in certain cases, including the above recreational games, optimal equilibria can be computed efficiently. In most other cases, we show that computing an optimal equilibrium is at least $\mathsf{NP}$-hard or $\mathsf{coNP}$-hard. Lastly, we experimentally validate our approach by computing exact equilibria for complete 5- and 6-player Avalon instances whose size in terms of number of information sets is larger than $10^{56}$.
@@ 1.1 Main Modeling Contributions
In defining communication extensions, we must bound the length of the communication, that is, how many rounds of communication occur in between every move of the game, and how many distinct messages can be sent on each round. To do this, we fix a finite message space [^4] Note that, if the message space is of size $M$, a message can be sent in $O(\log M)$ bits. of size $M$ and length of communication $R$, and in our definition of equilibrium we will take a supremum over $M$ and $R$. This will allow us to consider arbitrarily complex message spaces (i.e., $M$ and $R$ arbitrarily large) while still only analyzing finite games: for any fixed $M$ and $R$, the resulting game is a finite hidden-role game. We will show that our positive results (upper bounds) only require $\log M=R={\mathsf{polylog}}(|H|,1/\varepsilon)$, where $|H|$ is the number of nodes (histories) in the game tree and $\varepsilon$ is the desired precision of equilibrium.
@@ 1.2 Main Computational Contributions
If private communication is disallowed, the hidden-role value problem is $\mathsf{NP}$-hard. If the MIN-team is uncoordinated, the problem is $\mathsf{coNP}$-hard. If both, the problem is $\Sigma_{2}^{\mathsf{P}}$-hard. All hardness reductions hold even when the MIN-team is a minority and the number of players is an absolute constant.
@@ 1.4 Examples
Consider a $n$-player version of matching pennies (with $n>2$), which we denote as $\textsf{{MP}}(n)$. One player is chosen at random to be the adversary (team MIN). All $n$ players then simultaneously choose bits $b_{i}\in\{0,1\}$. Team MAX wins (gets utility $1$) if and only if all $n$ bits match; else, team MAX loses (gets utility 0).
With no communication, the value of this game is $1/2^{n-1}$: it is an equilibrium for everyone to play uniformly at random. Public communication does not help, because, conditioned on the public transcript, bits chosen by players must be mutually independent. Thus, the adversary can do the following: pretend to be on team MAX, wait for all communication to finish, and then play 0 if the string of all ones is more conditionally likely than the string of all $1$s, and vice-versa.
With private communication, however, the value becomes $1/(n+1)$. Intuitively, the MAX-team should attempt to guess who the MIN-player is, and then privately discuss among the remaining n1n-1 players what bit to play. We defer formal proofs of the above game values to Section 5, because they rely on results in Section 4.1.
@@ 2 Preliminaries
    A set of $n$ players, identified with the integers $n]=\{1,\dots,n\}$, and an extra player C denoting the chance player modeling stochasticity.
@@
    For each player $i\in[n]$, a utility function $u_{i}:Z\to[-1,1]$.
@@
and the TMEs are the strategy profiles $x$ that achieve the maximum value. Notice that the TME problem is nonconvex, since the objective function $u$ is nonlinear as a function of $x$ and $y$. As such, the minimax theorem does not apply, and swapping the teams may not preserve the solution. Computing an (approximate) TME is $\Sigma_{2}^{\mathsf{P}}$-complete in extensive-form games (Zhang et al., 2023a).
@@ 3 Equilibrium Concepts for Hidden-Role Games
    The utility of a player $i$ is defined completely by its team: there is a $u:Z\to\mathbb{R}$ for which $u_{i}(z)=u(z)$ if player $i$ is on team MAX at node $z$, and u(z)-u(z) otherwise. [^10] While at a first look this condition is similar to the one in ATGs, we remark that in this case the number of players in a team depends on the roles assigned at the start. The same considerations as Footnote 8 on the zero-sum rescaling of the utilities hold.
@@ 3.3 Equilibrium Notions
An $\varepsilon$-uncoordinated hidden-role equilibrium of $\Gamma$ with a particular notion of communication $c\in\{\textsf{none, pub, priv}\}$ is a tuple $(M,R,x)$ where $x$ is a MAX-strategy profile in $\textup{{USplit}}(\textup{{Comm}}_{c}^{M,R}(\Gamma))$ of value at least $\textsf{{UVal}}_{c}(\Gamma)-\varepsilon$. The $\varepsilon$-coordinated hidden-role equilibria is defined analogously, again with CSplit and CVal instead of USplit and UVal.
@@ 4.1 Computing Private-Communication Equilibria
Let $\Gamma$ be a hidden-role game with $k<n/2$. Then $\textsf{{CVal}}_{\textsf{{priv}}}(\Gamma)=\textsf{{CVal}}_{\textsf{{priv}}}(\Gamma^{*})$, where $\Gamma^{*}$ is $\Gamma$ with a mediator added, and moreover this value can be computed in $|H|^{O(k)}$ time by solving a zero-sum game of that size. Moreover, an $\varepsilon$-hidden-role equilibrium with private communication and $\log M=R={\mathsf{polylog}}(|H|,1/\varepsilon)$ can be computed and executed by the MAX-players in time ${\mathsf{poly}}(|H|^{k},\log(1/\varepsilon))$.
@@ 4.2 Computing No/Public-Communication Equilibria
In this section, we consider games with no communication or with public-communication and a coordinated minority. Conversely to the private-communication case of Section 4.1, in this case the problem of computing the value of a hidden-role equilibrium is in general $\mathsf{NP}$-hard.
@@
Even in 22-vs-11 games with public roles and $\varepsilon=1/{\mathsf{poly}}(|H|)$, computing the TME value (and hence also the hidden-role value, since adversarial team games are a special case of hidden-role games) with public communication is $\mathsf{NP}$-hard.
@@
Coordinated, Minority $\mathsf{NP}$}-complete (von Stengel and Koller, 1997) $\mathsf{NP}$-hard $\mathsf{P}$ [Thm. 4.3]
@@
Minority $\Sigma_{2}^{\mathsf{P}}$-complete $\Sigma_{2}^{\mathsf{P}}$-hard $\mathsf{coNP}$-hard
@@
Even with a constant number of players, a minority adversary team, and $\varepsilon=1/{\mathsf{poly}}(|H|)$, computing the uncoordinated value of a hidden-role game is $\mathsf{coNP}$-hard with private communication and $\Sigma_{2}^{\mathsf{P}}$-hard with public communication or no communication.
@@ 5 Worked Example
For the public-communication value, observe that, conditioned on the transcript, the bits chosen by the players must be mutually independent of each other. Thus, the adversary can do the following: pretend to be on team MAX, wait for all communication to finish, and then play 0 if the string of all ones is more conditionally likely than the string of all $1$s, and vice-versa [^13] In general, computing the conditional probabilities could take exponential time, but when defining the notion of value here, we are assuming that players have unbounded computational resources. This argument not work for computationally-bounded adversaries. Indeed, if the adversary were computationally bounded, MAX would be able to use cryptography to build private communication channels and thus implement a mediator, allowing our main positive result Theorem 4.3 to apply..
@@
(Lower bound) The players simulate a mediator using multi-party computation (see Theorems 4.3 and 4.2). Consider the following strategy for the mediator. Sample a string $b\in\{0,1\}^{n}$ uniformly at random from the set of $2n+2$ strings that has at most one mismatched bit. Recommend to each player $i$ that they play $b_{i}$.
@@
    Exactly one other bit of $b$ is 11 (n1n-1 ways).
@@ 6.1 The Price of Hidden Roles
Call a hidden-role game normalized if the value of $\textup{{PublicTeam}}(\Gamma)$ is $1$, that is, the MAX-team can win deterministically if the teams are publicly revealed. This assumption is typical in the example games that we consider; for example, it is easy to see that Avalon satisfies it.
@@
Let $D\in\Delta(\{{\color[rgb]{0.12109375,0.46484375,0.70703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.12109375,0.46484375,0.70703125}\textbf{{{MAX}}}},{\color[rgb]{0.12109375,0.46484375,0.70703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.12109375,0.46484375,0.70703125}\textbf{{{MAX}}}}\}^{n})$ be any distribution of teams assignments such that MAX is always a strict majority. Let $\mathcal{G}_{n,D}$ be the class of all normalized hidden-role games with $n$ players and team assignment distribution $D$. Then the price of hidden roles of $\mathcal{G}_{n,D}$ is exactly the largest probability assigned to any team by $D$, that is,
@@
(Upper bound) Team MAX assumes that the true MAX-team is exactly the team $t^{*}$. Then MAX gets utility at most a factor of $1/p^{*}$ worse than the TME value of $\textup{{PublicTeam}}(\Gamma)$: if the assumption is correct, then MAX gets the TME value; if the assumption is incorrect, MAX gets value at least 0 thanks to the condition on MAX’s utilities in Definition 6.2.
(Lower bound) Consider the following game $\Gamma$. Nature first selects a team assignment $t\sim D$ and each player privately observes its team assignment. Then, all players are simultaneously asked to announce what they believe the true team assignment is. The MAX-team wins if every MAX-player announces the true team assignment. If MAX wins, MAX gets utility $1$; otherwise MAX gets utility 0.
@@
Let $\mathcal{G}_{n,k}$ be the class of all normalized hidden-role games where the number of players and adversaries are always exactly $n$ and $k$ respectively, with $k<n/2$. The price of hidden roles in $\mathcal{G}_{n,k}$ is exactly $\binom{n}{k}$.
@@ 7 Experimental Evaluation: Avalon
No special roles (Resistance) 3 / 10 == 0.3000* 1 / 3 $\approx$ 0.3333*
Merlin 2 / 3 $\approx$ 0.6667* 3 / 4 == 0.7500*
@@
A game of Avalon proceeds, generically speaking, as follows. There are $n$ players, $\left\lceil n/3\right\rceil$ of which are randomly assigned to team MIN and the rest to team MAX. Team MIN is informed. Some special rules allow players observe further information; for example, Merlin is a MAX-player who observes the identity of the players on team MIN, except the MIN-player Mordred, and the MAX-player Percival knows Merlin and Morgana (who is on team MIN), but does not know which is which. Players are assigned roles by means of privately-dealt role cards. The role cards are treated as indistinguishable. [^14] In a typical Avalon card deck, the MAX “generic” (i.e., non-Merlin, non-Percival) role cards are actually distinguishable. As we have discussed in Section 1.4, this can have an impact on optimal play. However, it is a strong convention at all levels of human play that players do not exploit the distinguishability of role cards. We therefore generic MAX role cards as indistinguishable.
The game proceeds in five rounds. In each round, a leader publicly selects a certain number of people (defined as a function of the number of players and current round number) to go on a mission. Players then publicly vote on whether to accept the leader’s choice. If a strict majority vote to accept, the mission begins; otherwise, leadership goes to the player to the left. If four votes pass with no mission selected, there is no vote on the fifth mission (it automatically gets accepted). If a MIN-player is sent on a mission, they have the chance to fail the mission. The goal of MAX is to have three missions pass. If Merlin is present, MIN also wins by correctly guessing the identity of Merlin at the end of the game. Avalon is therefore parameterized by the number of players and the presence of the extra roles Merlin, Mordred, Percival, and Morgana.
@@
    A parallelized version of the PCFR+ algorithm (Farina et al., 2021), a scalable no-regret learning algorithm. PCFR+ was able to find an approximate equilibrium with exploitability $10^{-3}$ in less than 10 minutes in the largest game instance, and was able to complete 10,000 iterations in under two hours for each game.
@@ Acknowledgements
The work of Prof. Sandholm’s group is supported by the National Science Foundation under grants RI2312342 and RI-1901403, and by the ARO under award W911NF2210266. Nicola Gatti is supported by FAIR (Future Artificial Intelligence Research) project, funded by the NextGenerationEU program within the PNRR-PE-AI scheme (M4C2, Investment 1.3, Line on Artificial Intelligence). We thank Justin Raizes for useful discussions on multi-party computation, and Cristian Palma Foster for pointing out an error in an earlier version of the statement of Theorem 6.3.
@@ B.1 Secure MPC
In secure MPC, there is a (possibly randomized) function $f:X^{n}\to X^{n}$ defined by a circuit with $N$ nodes. A subset $K\subset[n]$ of size <n/2<n/2 has been corrupted. Each honest player $i\in[n]\setminus K$ holds an input $x_{i}\in X$. The goal is to design a randomized messaging protocol, with private communication, such that, regardless of what the corrupted players do, there exist inputs $\{x_{j}:j\in K\}$ such that:
@@
    (Privacy) No subset of <n/2<n/2 players can learn any information except their own output $y_{i}$. That is, the players cannot infer any extra information from analyzing the transcripts of the message protocol than what they already know.
    This can be intuitively modeled as follows. If any minority of colluded players were to analyze the empirical distributions of transcripts of the protocol for any input-output tuple, then such distribution would be fully explained in terms of their input-outputs, leaving no conditional dependence on other players’ input-outputs. Formally, for any such subset $K\subseteq[n]$ of size <n/2<n/2, there exists a randomized algorithm ${\sf Sim}_{K}$ that takes the inputs and outputs of the players in set $K$, and reconstructs transcripts $T$, such that for all $x\in X^{n}$, we have
[…]
@@ B.2 Verifiable Secret Sharing
In the verifiable secret sharing (VSS) problem, the goal is to design a function ${\sf Share}:X\to{\bar{X}}^{n}$ , where ${\bar{X}}=\{0,1\}^{{\mathsf{poly}}(\ell,\lambda,n)}$, that shares a secret $x\in X$ by privately informing each player $i$ of its piece ${\sf Share}(x)_{i}$, such that:
@@
    (Reconstructibility) Any subset of >n/2>n/2 players can recover the secret fully, even if the remaining players are adversarial. That is, there exists a function ${\sf Reconstruct}:X^{n}\to X$ such that, where $(x_{1},\dots,x_{n})\leftarrow{\sf Share}(x)$, we have ${\sf Reconstruct}(x_{1}^{\prime},\dots,x_{n}^{\prime})=x$ so long as $x_{i}^{\prime}=x_{i}$ for >n/2>n/2 players $i$.
@@
    (Privacy) No subset of <n/2<n/2 players can learn any information about the secret $x$. That is, for any such subset $K$ and any secret $x$, there exists a distribution ${\sf Sim}_{K}\in\Delta(X^{n})$ such that
@@
There exist algorithms Share and Reconstruct, with runtime ${\mathsf{poly}}(\ell,\lambda,n)$, that implement robust secret sharing.